OpenStudy (anonymous):

In the 2000 season, quarterback Jeff Blake of the New Orleans Saints had 9 passes intercepted out of 302 attempts. What percent of Jeff Blake’s passes were intercepted? Round to the nearest tenth if necessary.

hero (hero):

9/302 = x/100 302x = 900 x = 900/302 x is approximately 3

hero (hero):

so 3%